# Behind the Magic, MERLIM: Multi-modal Evaluation Benchmark for Large Image-Language Models

### v2

@@ 3.1 Object Recognition
We identify all the nouns in $r$ and compare them against the ground truth labels of MS-COCO. To do this, we split the sentence by words and determine if it is a noun or not using the spaCy library (Honnibal & Montani, 2017). We discard words that refer to places or areas like room, street, square, etc. We consider compound nouns as one, for instance, ”the back of the car” and ”ice cream”. We designate a prediction as a True Positive if the predicted word matches or is a synonym of a ground truth element. Likewise, a prediction that does not match any ground=truth label or any of its synonyms is labeled as a False Positive. Finally, any ground truth class that can not be found in the prediction (including their synonyms) is regarded as a False Negative. These definitions enable us to calculate precision, recall, and F1 metrics for all IT-LVLMs.
@@ 3.2 Inter-object Relationship Understanding
It is challenging to direct the IT-LVLM attention to a specific instance when there are multiple objects of the same class. This makes the visual grounding assessment ambiguous, as the target relationship might randomly exist among two other entities with the same semantic classes. We mitigate this ambiguity by substituting nouns that have two or more instances in the image with another class noun. We follow two strategies to select the new object categories: (1) We randomly sample another object category from the ground truth classes of MS-COCO which are not included in the image ground truth. We name this set the Random Set. (2) We ask the ChatGPT API (OpenAI, 2023) to select an object category to generate a meaningful relationship (in terms of language). This set of relationships is named the Curated Set. The expected answer to the queries in any of these sets are always negative, as the relationship no longer has a valid visual grounding.
@@ 3.3 Object Counting
Finally, we also evaluate the counting capabilities of IT-LVLMs. Although this task is common in VQA benchmarks, even state-of-the-art methods struggle to provide accurate results. In particular, the counting task requires the models to have strong object recognition and instance detection capabilities that enable them to count in a discrete fashion (Trott et al., 2018; Zhang et al., 2018).
@@ 4 MERLIM Results
We proceed with the experimental assessment of several representative state-of-the-art Instruction Tuning Large Visual Language Models, following the protocol outlined in MERLIM. Specifically, we evaluate BLIP-2 (with FlanT5xl LLM) (Li et al., 2023a), InstructBLIP (with Vicuna-7B v1.1, Vicuna-13B v1.1 and FlanT5xl LLMs) (Dai et al., 2023), MiniGPT-4 (with Vicuna-7B v0 and Vicuna-13B v0 LLMs) (Zhu et al., 2023), LLaVA-1.5 (with Vicuna-7B v1.5 LLM) (Liu et al., 2023), xGen-MM (with Phi-3 Mini 3.8B LLM) (Research, 2024), Kosmos-2 (with MAGNETO LLM) (Peng et al., 2024), InternLM-XComposer2-VL (with InternLM-7B LLM) (Dong et al., 2024), and Qwen-VL-Chat (with Qwen-7B LLM) (Bai et al., 2023). We outline and analyze the results across the three evaluation tasks in MERLIM.
@@ 4.1 Object Recognition
Figure 2 outlines the performance (F1 score) of the IT-LVLMs in MERLIM for each prompt. Interestingly, while ChatGPT (OpenAI, 2023) and Vicuna-33B (Chiang et al., 2023) consider the prompts semantically identical, only BLIP-2 (Li et al., 2023a), and Kosmos-2 (Peng et al., 2024) exhibit consistent performance across all five instructions. These models exhibit low standard deviations of 0.59 and 0.67, respectively. Notably, BLIP-2 achieves the best average performance with 40.55%, even outperforming its successor, InstructBLIP. This superior consistency of BLIP-2 over InstructBLIP might stem from it lacking an instruction-aware Q-former module.
@@
Table 2: Results for Inter-object Relationship Understanding. We report the Accuracy (Acc) of the IT-LVLMs in the two sets: one with randomly sampled relationships (Random Set) and another where the relationships are curated by an LLM (Curated Set). We compute the Accuracy with the affirmative ($\bf Acc$) and negative ($\bf Acc^{neg}$) versions of the relationships to validate the models’ answers. We also present the Acc in both sets of COCO images, the edited ($\bf Acc_{ed}$) and Orginal ($\bf Acc_{org}$), and the absolute accuracy difference $\bf\Delta Acc=Abs(Acc_{org}-Acc_{ed})$. The Curated set constitutes a more challenging set for all methods because. No IT-LVLM is superior across all evaluated scenarios.
@@
MERLIM can break down Hallucination events into regular and hidden Hallucinations. To this end, we compare the precision of IT-LVLMs on both the original and edited image sets. Figure 3 presents results for two of the five designed prompts [^1] We chose Questions 1 and 2 as they are, on average, the best-performing prompts by F1 Score. Regarding regular hallucinations, 5 of the 11 models we tested (Dai et al., 2023; Dong et al., 2024; Li et al., 2023a; Liu et al., 2023; Research, 2024) achieved a precision near or above 50% on the original image set in at least one of the selected questions. Surprisingly, these models exhibit the largest performance gap on the edited image set, with an average drop of 8.05%. This empirical evidence indicates that the best performing methods are also the ones which suffer the most from hidden hallucinations.
Some replies lack direct visual grounding (as the corresponding visual instance was removed from the image in the inpainting process), however these objects persist in the language prediction. This result suggests that the best performing IT-LVLMs are obtaining an advantage by prioritizing the global visual context or previously predicted language tokens over the effective visual grounding. We provide further evidence to support this observation in Table 5 (Supplementary Material). In this table, we show that the gradient contributions of the previously predicted language tokens are, on average, larger than those of the image. This suggest that top-performing models are improving their performance by hallucinating objects which are largely determined by the initial answer tokens.
@@
Figure 4: Results on Object Counting task. Subfigure a) reports the mean absolute error of IT-LVLMs in the original and edited sets. LlaVa v1.5 (Liu et al., 2023) demonstrates better performance across both sets. On average, BLIP2 and MiniGPT make the worst estimations. Every model reduces its performance in the edited set, indicating a weak correlation between visual grounding and model response. Subfigure b plots the accuracy for the yes/no questions designed to validate answer consistency for the main counting task. LlaVa exhibits remarkable consistency across both sets, significantly outperforming other methods. Other IT-LVLMs perform close to random chance, suggesting a bias towards the question format in their answers.
@@ 4.2 Inter-object Relationship Understanding
Table 2 shows that all tested IT-LVLMs exhibit an average performance drop of 7.49% when transitioning from the original to the edited set. This performance gap, is nearly the same gap in the object recognition task, suggesting that IT-LVLMs may also be hallucinating certain visual relationships due to the global visual context and the inherent biases within the language models. MiniGPT-4 (Zhu et al., 2023) (with Vicuna-7B or 13B) attains the lowest accuracy overall, although it compensates with the smallest variability between the original and edited sets ($\bf\Delta Acc$). With MiniGPT-4’s LLM and Visual stream being comparable in size to other models, the reduced performance could be attributed to a different Vicuna version (V0 for MiniGPT-4) and a relatively small fusion module ($\theta_{h}$), consisting of a single linear layer. It is noteworthy that no IT-LVLM consistently outperforms the others in all the evaluated scenarios. However, Qwen-VL-Chat (Bai et al., 2023) exhibits the best performance for affirmative relationships in both original and edited images, regardless of the set (random or curated).
@@
The curated set requires the IT-LVLMs to build a better understanding of visual information to reply about the visual relationships. In short, relationships can not be trivially resolved from the language query (see Section 3.2). As summarized in Table 2, all the evaluated IT-LVLMs significantly reduce their performance in the edited set by 11.63% when compared to the curated set. Furthermore, the average performance gap between the original and edited sets ($\bf\Delta Acc$) has risen to 16.15% more than double the gap observed in the random set.
@@ 4.3 Object Counting
In Figure 4, we report the mean absolute error (MAE) and its standard deviation for all tested IT-LVLMs on both the edited (blue) and original (orange) in MERLIM. All IT-LVLMs exhibit a significant increase in error on the edited set. This indicates that the intrinsic error of the models in the counting task is not the only source of error but also a lack of visual grounding in their responses, which leads to hallucinations.
@@
To further investigate language biases, we also compute the accuracy metric for the yes/no questions, which validate the answers given by the IT-LVLM in the initial prediction (i.e. Are there N objects?). We highlight that all methods still perform better in the original image set, especially the largest IT-LVLMs, like InstructBLIP over Vicuna-13B. However, the drop in performance remains significant in the edited set. LLaVA (Liu et al., 2023) exhibits the highest consistency, with over 80% of the replies confirming the initial assessment (regardless if the count was accurate or not) for both the edited and original sets. Similarly, the LLaVA IT-LVLM also has the smallest average error and the smallest standard deviation on the initial counting, clearly outperforming every other IT-LVLM in this particular task.
@@ 5 Discussion & Conclusions
Unlike previous benchmarks, MERLIM stands out for its ability to define and study a novel type of hallucination: hidden hallucinations. These are more challenging to detect because they are seemingly correct predictions but in fact lack a meaningful visual grounding. Our findings suggest that top-performing models might achieve higher scores (in part) due to these hidden hallucinations. Moreover, our analysis also identifies potential causes for this phenomenon: language biases and a lack of robust visual grounding.
@@
MERLIM comprises a total of 300664 unique image-question pairs across all tasks, which may limit its usability for evaluating and prototyping new models. Therefore, we recommend using the subset of MERLIM designed for the Object Recognition Task, where image inpainting removes an entire category. This subset, used in Figure 3, contains 5608 edited and 3037 original images.
@@ Conclusion.
Appendix A Supplementary Material
@@
A.1 Analysis of Inpainted Images
To analyze the visual changes in the images after the inpainting process, we first assess that the selected inpainting strategy induces a minimal change in the global image features and that the inpainted masks effectively hide the object instance from object detectors. We use YOLOv7 (Wang et al., 2023) to verify the absence of the object in the edited image and ResNet50 (He et al., 2016) to evaluate the global feature similarity with the original image. For each image in MERLIM, we enforce two essential criteria: (1) YOLOv7 must not generate a detection box with an Intersection over Union (IoU) greater than 0.7 relative to the ground-truth bounding box of the removed instance, and (2) the cosine similarity between the ResNet50 features of the edited image and the original image must exceed 0.7.
For a more in-depth assessment of the visual changes in the inpainted images, we use another box detector, Mask R-CNN (He et al., 2017), to verify if each predicted box in the inpainted image has a corresponding box with the same class and a similar spatial location (i.e., high Intersection over Union) in the original image.
We perform this analysis for bounding boxes with confidence scores above 0.7 and 0.5, summarizing the results in Table 3.
@@ A.1 Analysis of Inpainted Images
Out of a total of 762782 predicted boxes in the inpainted image set, we found that only 4455 boxes (about 0.59%) could not be mapped to the boxes detected in the original image with the same class and an overlap of at least 0.5 IoU. In comparison, 726969 boxes (96.8%) in the inpainted images have a matching box with the same class and a nearly identical spatial location (IoU ¿ 0.7) in the original image. Despite the inpainting strategy being relatively under-explored, we conclude that the observed performance gaps in MERLIM should not be attributed to the minimal discrepancies identified by representative image networks (ResNet50, YOLOv7, Mask R-CNN) but rather to hallucination events in the IT-LVLM.
Figure 5: Number of nouns predicted. We extract the nouns using the spaCy library from the answers of the models from all the formulated prompts on the original image set. It is worth noting that BLIP-2 predicts a consistent number of nouns across all the prompts, unlike the other methods.
Figure 6: Precision and Recall on the original image set.. We compare the Precision and the Recall of IT-LVLMs on the original image set across the five proposed prompts $P$ in the sub-figure a) and b), respectively. It is worth noting that the biggest models, such as InstructBLIP and MiniGPT-4 with Vicuna13B and LLaVa, get the highest recall and lowest precision.
A.2 Object Hallucination
In Figure 5, we provide further evidence of instruction bias in IT-LVLMs. Despite the semantic equivalence of the proposed prompts, all methods, except for BLIP-2 (Li et al., 2023a) and Kosmos-2 (Peng et al., 2024), generate varying numbers of nouns for each prompt. Figure 6 illustrates that while BLIP-2 produces the shortest answers (in terms of noun count), it consistently achieves higher precision compared to other methods, albeit with lower recall. This indicates that BLIP-2 is less prone to hallucination, likely because it was trained to generate concise captions rather than detailed descriptions. On the other hand, larger models trained with instructional data, such as InstructBLIP and MiniGPT-4 with Vicuna-13B (Dai et al., 2023; Zhu et al., 2023), LLaVa (Liu et al., 2023), and xGen-MM (Research, 2024), tend to generate longer answers on average. This increases the likelihood of hallucinations, particularly when visual grounding is insufficient, resulting in higher recall but lower precision compared to BLIP-2.
A.3 LLMs vs IT-LVLMs

### v3

@@ 3.1 Object Recognition
We identify all the nouns in $r$ and compare them against the ground truth labels of MS-COCO. To do this, we split the sentence by words and determine if it is a noun or not using the spaCy library [21]. We discard words that refer to places or areas like room, street, square, etc. We consider compound nouns as one, for instance, "the back of the car" and "ice cream". We designate a prediction as a True Positive if the predicted word matches or is a synonym of a ground truth element. Likewise, a prediction that does not match any ground truth label or any of its synonyms is labeled as a False Positive. Finally, any ground truth class that can not be found in the prediction (including their synonyms) is regarded as a False Negative. These definitions enable us to calculate precision, recall, and F1 metrics for all IT-LVLMs.
@@ 3.2 Inter-object Relationship Understanding
It is challenging to direct the IT-LVLM attention to a specific instance when there are multiple objects of the same class. This makes the visual grounding assessment ambiguous, as the target relationship might randomly exist among two other entities with the same semantic classes. We mitigate this ambiguity by substituting nouns that have two or more instances in the image with another class noun. We follow two strategies to select the new object categories: (1) We randomly sample another object category from the ground truth classes of MS-COCO which are not included in the image ground truth. We name this set the Random Set. (2) We ask the ChatGPT API [41] to select an object category to generate a meaningful relationship (in terms of language). This set of relationships is named the Curated Set. The expected answer to the queries in any of these sets are always negative, as the relationship no longer has a valid visual grounding.
@@ 3.3 Object Counting
Finally, we also evaluate the counting capabilities of IT-LVLMs. Although this task is common in VQA benchmarks, even state-of-the-art methods struggle to provide accurate results. In particular, the counting task requires the models to have strong object recognition and instance detection capabilities that enable them to count in a discrete fashion [48, 55].
@@ 4 MERLIM Results
We proceed with the experimental assessment of several representative state-of-the-art Instruction Tuning Large Visual Language Models, following the protocol outlined in MERLIM. Specifically, we evaluate BLIP-2 (with FlanT5xl LLM) [30], InstructBLIP (with Vicuna-7B v1.1, Vicuna-13B v1.1 and FlanT5xl LLMs) [9], MiniGPT-4 (with Vicuna-7B v0 and Vicuna-13B v0 LLMs) [58], LLaVA-1.5 (with Vicuna-7B v1.5 LLM) [36], xGen-MM (with Phi-3 Mini 3.8B LLM) [52], Kosmos-2 (with MAGNETO LLM) [43], InternLM-XComposer2-VL (with InternLM-7B LLM) [11], and Qwen-VL-Chat (with Qwen-7B LLM) [4]. Following, We outline and analyze the results across the three evaluation tasks in MERLIM.
@@ 4.1 Object Recognition
Figure 2 outlines the performance (F1 score) of the IT-LVLMs in MERLIM for each prompt. Interestingly, while ChatGPT [41] and Vicuna-33B [8] consider the prompts semantically identical, only BLIP-2 [30], and Kosmos-2 [43] exhibit consistent performance across all five instructions. These models boast low standard deviations of 0.59 and 0.67, respectively. Notably, BLIP-2 achieves the best average performance with 40.55%, even outperforming its successor, InstructBLIP. This superior consistency of BLIP-2 over InstructBLIP might stem from its lacking an instruction-aware Q-former module. This module’s presence in InstructBLIP could make it more susceptible to slight variations in the phrasing of instructions.
@@
Table 2: Results for Inter-object Relationship Understanding. We report the Accuracy (Acc) of the IT-LVLMs in the two sets: one with randomly sampled relationships (Random Set) and another where the relationships are curated by an LLM (Curated Set). We compute the Accuracy with the affirmative ($\bf Acc$) and negative ($\bf Acc^{neg}$) versions of the relationships to validate the models’ answers. We also present the Acc in both sets of COCO images, the edited ($\bf Acc_{ed}$) and Original ($\bf Acc_{org}$), and the absolute accuracy difference $\bf\Delta Acc=Abs(Acc_{org}-Acc_{ed})$. The Curated set constitutes a more challenging set for all methods because. No IT-LVLM is superior across all evaluated scenarios.
@@
MERLIM can break down Hallucination events into regular and hidden Hallucinations. To this end, we compare the precision of IT-LVLMs on both the original and edited image sets. Figure 3 presents results for two of the five designed prompts [^1] We chose Questions 1 and 2 as they are, on average, the best-performing prompts by F1 Score. Regarding regular hallucinations, 5 of the 11 models we tested InstructBLIP [9] InternLM-XComposer2-VL [11], BLIP-2 [30], LLaVA-1.5 [36] and xGen-MM [52] achieved a precision near or above 50% on the original image set in at least one of the selected questions. Surprisingly, these models exhibit the largest performance gap on the edited image set, with an average drop of 8.05%. This empirical evidence indicates that the best performing methods are also the ones which suffer the most from hidden hallucinations.
Clearly some replies lack direct visual grounding (as the corresponding visual instance was removed from the image in the inpainting process), however these objects persist in the language prediction. This result suggests that the best performing IT-LVLMs are obtaining an advantage by prioritizing the global visual context or previously predicted language tokens over the effective visual grounding. We provide further evidence to support this observation in the Supplementary Material. In this table, we show that the gradient contributions of the previously predicted language tokens are, on average, larger than those of the image. This suggests that top-performing models are improving their performance by hallucinating objects, which are largely determined by the initial answer tokens.
@@ 4.2 Inter-object Relationship Understanding
Table 2 shows that all tested IT-LVLMs exhibit an average performance drop of 7.49% when transitioning from the original to the edited set. This performance gap, is nearly the same gap in the object recognition task, suggesting that IT-LVLMs may also be hallucinating certain visual relationships due to the global visual context and the inherent biases within the language models. MiniGPT-4 [58] (with Vicuna-7B or 13B) attains the lowest accuracy overall, although it compensates with the smallest variability between the original and edited sets ($\bf\Delta Acc$). Despite MiniGPT-4’s LLM and Visual stream being comparable in size to other models, the reduced performance could be attributed to a different Vicuna version (V0 for MiniGPT-4) and a relatively small fusion module ($\theta_{h}$), consisting of a single linear layer. It is noteworthy that no IT-LVLM consistently outperforms the others in all the evaluated scenarios. However, Qwen-VL-Chat [4] exhibits the best performance for affirmative relationships in both original and edited images, regardless of the set (random or curated).
Figure 4: Results on Object Counting task. Subfigure a) reports the mean absolute error of IT-LVLMs in the original and edited sets. LlaVa v1.5 [36] demonstrates better performance across both sets. On average, BLIP2 and MiniGPT make the worst estimations. Every model reduces its performance in the edited set, indicating a weak correlation between visual grounding and model response. Subfigure b plots the accuracy for the yes/no questions designed to validate answer consistency for the main counting task. LlaVa exhibits remarkable consistency across both sets, significantly outperforming other methods. Other IT-LVLMs perform close to random chance, suggesting a bias towards the question format in their answers.
@@
The curated set requires the IT-LVLMs to better understand visual information to reply about the visual relationships. In short, relationships can not be trivially resolved from the language query (see Section 3.2). As summarized in Table 2, all the evaluated IT-LVLMs significantly reduce their performance in the edited set by 11.63% when compared to the curated set. Furthermore, the average performance gap between the original and edited sets ($\bf\Delta Acc$) has risen to 16.15% more than double the gap observed in the random set.
@@ 4.3 Object Counting
In Figure 4, we report the mean absolute error (MAE) and its standard deviation for all tested IT-LVLMs on both the edited (blue) and original (orange) in MERLIM. All IT-LVLMs exhibit a significant increase in error on the edited set. This indicates that the models’ intrinsic error in the counting task is not correlated with the visual grounding, but rather a lack of visual grounding in their responses, leading to hallucinations.
@@
To further investigate language biases, we also compute the accuracy metric for the yes/no questions, which validate the answers given by the IT-LVLM in the initial prediction (i.e. Are there N objects?). We highlight that all methods still perform better in the original image set, especially the largest IT-LVLMs, like InstructBLIP over Vicuna-13B. However, the drop in performance remains significant in the edited set. LLaVA [36] exhibits the highest consistency, with over 80% of the replies confirming the initial assessment (regardless if the count was accurate or not) for both the edited and original sets. Similarly, the LLaVA IT-LVLM also has the smallest average error and the smallest standard deviation on the initial counting, clearly outperforming every other IT-LVLM in this particular task.
@@ 5 Discussion & Conclusions
Unlike previous benchmarks, MERLIM stands out for its ability defining and studying a novel type of hallucination: hidden hallucinations. These are more challenging to detect because they are seemingly correct predictions but in fact lack a meaningful visual grounding. Our findings suggest that top-performing models might achieve higher scores (in part) due to these hidden hallucinations. Moreover, our analysis also identifies potential causes for this phenomenon: language biases and a lack of robust and detailed visual grounding.
@@
MERLIM comprises a total of 300,664 unique image-question pairs across all tasks, which may limit its usability for evaluating and prototyping new models. Therefore, we recommend using the subset of MERLIM designed for the Object Recognition Task, where image inpainting removes an entire category. This subset, used in Figure 3, contains 5608 edited and 3037 original images.
@@ Conclusion.
Acknowledgments.
This work is supported by the KAUST Center of Excellence for Generative AI under award number 5940. The computational resources are provided by IBEX, which is managed by the Supercomputing Core Laboratory at KAUST.
Supplementary Material
@@ Acknowledgments.
6 Analysis of Inpainted Images
To analyze the visual changes in the images after the inpainting process, we first assess that the selected inpainting strategy induces a minimal change in the global image features and that the inpainted masks effectively hide the object instance from object detectors. We use YOLOv7 [50] to verify the absence of the object in the edited image and ResNet50 [19] to evaluate the global feature similarity with the original image. For each image in MERLIM, we enforce two essential criteria: (1) YOLOv7 must not generate a detection box with an Intersection over Union (IoU) greater than 0.7 relative to the ground-truth bounding box of the removed instance, and (2) the cosine similarity between the ResNet50 features of the edited image and the original image must exceed 0.7.
For a more in-depth assessment of the visual changes in the inpainted images, we use another box detector, Mask R-CNN [20], to verify if each predicted box in the inpainted image has a corresponding box with the same class and a similar spatial location (i.e., high Intersection over Union) in the original image.
We perform this analysis for bounding boxes with confidence scores above 0.7 and 0.5, summarizing the results in Table 3(b).
@@ 6 Analysis of Inpainted Images
Out of a total of 762782 predicted boxes in the inpainted image set, we found that only 4455 boxes (about 0.59%) could not be mapped to the boxes detected in the original image with the same class and an overlap of at least 0.5 IoU. In comparison, 726969 boxes (96.8%) in the inpainted images have a matching box with the same class and a nearly identical spatial location (IoU $>$ 0.7) in the original image. Moreover, in Figure 5, we also provide visual examples of the resulting images after the inpainting process. As can be seen, the inpainting images are highly similar to the original ones. The editions are visually indistinguishable even when large objects are removed. Despite the inpainting strategy being relatively under-explored, we conclude that the observed performance gaps in MERLIM should not be attributed to the minimal discrepancies identified by representative image networks (ResNet50, YOLOv7, Mask R-CNN) but rather to hallucination events in the IT-LVLM.
Figure 5: Examples of Inpainting images. We present eight examples demonstrating the results of removing an object from the original image using the inpainting model proposed by [31]. As illustrated, the edits are visually seamless, even when large objects are removed, making them nearly indistinguishable from the original.
Figure 6: Number of nouns predicted. We extract the nouns using the spaCy library from the answers of the models from all the formulated prompts on the original image set. It is worth noting that BLIP2 predicts a consistent number of nouns across all the prompts, unlike the other methods.
Figure 7: Precision and Recall on the original image set.. We compare the Precision and the Recall of IT-LVLMs on the original image set across the five proposed prompts $P$ in the sub-figure a) and b), respectively. It is worth noting that the biggest models, such as InstructBLIP and MiniGPT4 with Vicuna13B and LlaVa, get the highest recall and lowest precision.
7 Object Hallucination
In Figure 6, we provide further evidence of instruction bias in IT-LVLMs. Despite the semantic equivalence of the proposed prompts, all methods, except for BLIP-2 [30] and Kosmos-2 [43], generate varying numbers of nouns for each prompt. Figure 7 illustrates that while BLIP-2 produces the shortest answers (in terms of noun count), it consistently achieves higher precision compared to other methods, albeit with lower recall. This indicates that BLIP-2 is less prone to hallucination, likely because it was trained to generate concise captions rather than detailed descriptions. On the other hand, larger models trained with instructional data, such as InstructBLIP and MiniGPT-4 with Vicuna-13B [9, 58], LLaVa [36], and xGen-MM [52], tend to generate longer answers on average. This increases the likelihood of hallucinations, particularly when visual grounding is insufficient, resulting in higher recall but lower precision compared to BLIP-2. As shown in Figure 8 shows that even a proprietary model such as GPT-4o-mini [41] presents a high hallucination rate.
Figure 8: Comparing GPT-4o-mini with the open-source models on the original and edited sets. (a) We compare the precision of IT-LVLMs on the edited and original image sets using one prompt. To analyze the hidden hallucination problem, we focus on the subset where the image inpainting removes an entire category. Notably, all methods lose performance on the edited image set. (b) We also report the F1 Score on the original and edited images using the same question.
8 LLMs vs IT-LVLMs